OpenStudy (horsegal244):

Rob measured the height of an art frame as 5.12 feet, but the actual height was 6 feet. What is the percent of error in Rob's measurement?

OpenStudy (horsegal244):

@Michele_Laino

OpenStudy (michele_laino):

the percent error is given by the subsequent computation: \[\frac{{6 - 5.12}}{{5.12}} \times 100 = ...\]

OpenStudy (horsegal244):

ok i got the answer ty